OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos2x

zepdrix (zepdrix):

Is that a square on the right side or a (2x) inside of cosine?

OpenStudy (anonymous):

@zepdrix, probably a \(\cos^2x\). Expanding the left side gives a variant of \(\cos^2x-\sin^2x\).

OpenStudy (anonymous):

@maywadee313, \[\sin x\left(\tan x\cos x-\cot x\cos x\right)=\cdots\\ \sin x\left(\frac{\sin x}{\cos x}\cos x-\frac{\cos x}{\sin x}\cos x\right)=\cdots\\ \sin x\left(\sin x-\frac{\cos^2 x}{\sin x}\right)=\cdots\\ \sin^2 x-\cos^2 x=\cdots \] Use the double angle identity for cosine for the next few steps: \[\cos2x=\cos^2x-\sin^2x\\~~~~~~~~~=2\cos^2x-1\]

OpenStudy (anonymous):

@SithsAndGiggles , whats the double angle identity?

OpenStudy (anonymous):

It's the last thing I listed: \(\cos2x=\cos^2x-\sin^2x\).

OpenStudy (anonymous):

but how do I use it? I never learned that

OpenStudy (anonymous):

Hey, you know what? Forget I mentioned "double angle" identity. You don't need to know about it for this specific problem. Just getting ahead of myself :) If you haven't learned it yet, you will eventually. Do you understand how I got to \(\sin^2x-\cos^2x\) in the last step?

OpenStudy (anonymous):

kind of. mind explaining?

OpenStudy (anonymous):

From first to second line, I simply used the definitions of \(\tan x\) and \(\cot x\): \[\tan x=\frac{\sin x}{\cos x}~~~\text{and}~~~\cot x=\frac{1}{\tan x}=\frac{\cos x}{\sin x}\] From second to third, I simplified the quantity inside the parentheses. From third to fourth, I distributed the sine on the outside.

OpenStudy (anonymous):

alright I understand that

OpenStudy (anonymous):

Okay, so the last few steps require using the Pythagorean identity: \(\sin^2x+\cos^2x=1\). You're familiar with that one, right? \[\sin^2x-\cos^2x=\left(1-\cos^2x\right)-\cos^2x\]

OpenStudy (anonymous):

yes. I am familiar with that for the most part.

OpenStudy (anonymous):

Okay, so the last step is to just combine like terms. Left side matches up with right, so question answered: \[1-\cos^2x-\cos^2x=1-2\cos^2x\]

OpenStudy (anonymous):

ohhh okay :) thanks !

OpenStudy (anonymous):

You're welcome